A reverse osmosis unit is used to obtain drinkable
Answer #1

given

weight per liter = 0.750 g of NaCl

molecular weight of NaCl = 58.5 g/ mol

molarity = 0.750/ 58.5 = 0.0128 M

osmotic pressure = i * M * R * T

i = vant hoff factor ( for NaCl 1.8)

M = Molarity mol/L

R = gas Constant (L atm / mol K ) = 0.08206

T= temperature = 298 K

osmotic pressure = 1.8 * 0.0128 * 0.0826 * 298

osmotic pressure = 0.564 atm = 428.64 torr